OpenStudy (anonymous):

An automobile purchased for use by the manager of a firm at a price of $24,000 is to be depreciated using the straight-line method over 5 years. What will be the book value of the automobile at the end of 3 years? (Assume the scrap value is $0.)

OpenStudy (anonymous):

i got p=-.5x+100

OpenStudy (ash2326):

how did you find this?

OpenStudy (anonymous):

i got 9600

OpenStudy (anonymous):

answer is 9600

OpenStudy (anonymous):

i'm going to post a new problem could you take a look at it?

OpenStudy (ash2326):

Yeah, correct:) could you show me your working?

OpenStudy (anonymous):

no i got it from someone else

OpenStudy (ash2326):

@zackwashere just don't aim to find answers, try to understand the concept behind it. I'll explain this problem

OpenStudy (ash2326):

We need to find the annual depreciation value of the automobile: Let it be d \[d=\frac{\text{cost}-\text{scrap value}}{\text{no. of years}}\] \(d=\frac{24000-0}{5}\) \(= 4800\) so each year its price will depreciate by 4800 after 3 years, it'll get reduce by 3*4800=14400 so the book value will be =>24000-14400=9600 do you understand this?